OpenStudy (jjuden):

Find the roots of the polynomial equation. x^2 – 8x – 48 = 0

OpenStudy (jjuden):

–6, 8 –4, 12 –12, 4 –8, 6

OpenStudy (anonymous):

x² - 8x - 48 = 0 x² - 4x = 48 + 4² x² - 4x = 48 + 16 (x - 4)² = 64 x - 4 = 8 x = 8 + 4, x = 12 x = - 8 + 4, x = - 4